OpenStudy (anonymous):

given a circle whose equation is x^2 +8x+y^2-4y=5 determine the equation of the circle that is translated to the right 4 units and up 2 units that has an area which is twice that of the given circle

OpenStudy (anonymous):

this may actually go under geometry, but it was one of my precal questions so I'm not sure

OpenStudy (jdoe0001):

so pretty much to get the Area of any circle all you need is the "radius" so in short, once you get the radius of the GIVEN circle, you'd get its Area double that Area, will equal \(\bf \large \pi r^2\) in the other circle, solve for "r" to get the radius of it once you have its radius, you have the equation, notice that (h, k) are given

OpenStudy (jdoe0001):

"translated to the right 4 units and up 2 units" <---- (h, k)

OpenStudy (jdoe0001):

so, all you really need is the radius

OpenStudy (jdoe0001):

"translated to the right 4 units and up 2 units" => \(\bf (x+4)^2+(y-2)^2 = r^2\)

OpenStudy (anonymous):

but how do I find the radius of the given circle? do i have to graph it??

OpenStudy (jdoe0001):

you complete the square, to get the equation

OpenStudy (anonymous):

of the origional equation or the one you just wrote?

OpenStudy (jdoe0001):

\(\bf x^2 +8x+y^2=5\)

OpenStudy (anonymous):

ok hold on i got this

OpenStudy (anonymous):

did you drop the -4y on purpose?

OpenStudy (jdoe0001):

"given a circle whose equation is x^2 +8x+y^2=5 determ...." <--- is the posting

OpenStudy (anonymous):

shoot! sorry let me fix that

OpenStudy (anonymous):

ok no sorry. its x^2 +8x+y^2-4y=5

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

(x+4)^2 (y-2)^2=25

OpenStudy (jdoe0001):

yes, that'd be the circle, and its Area = \(\bf 2 \pi r^2\)

OpenStudy (anonymous):

r = 25?

OpenStudy (jdoe0001):

\(\bf r^2 = 25 \implies r = 5\)

OpenStudy (anonymous):

50 pi

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

\(\bf r = 5\qquad \qquad Area = 2 \pi r^2 \implies Area = 50\pi\\ \textit{area of the other circle is double that } 100\pi \textit{thus }\\ 100\pi = \pi r^2 \implies \sqrt{\cfrac{100\cancel{\pi}}{\cancel{\pi}} }= r\)

OpenStudy (jdoe0001):

"translated to the right 4 units and up 2 units" => \(\bf (x+4)^2+(y-2)^2 = r^2 \implies (x+4)^2+(y-2)^2 = 10^2\)

OpenStudy (anonymous):

thank you so much! if i could give you another medal I would!

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

hmnmmm

OpenStudy (jdoe0001):

one sec ... .... I have.. one thing amiss

OpenStudy (jdoe0001):

I used above \(\bf 2\pi r^2 \) for the area which is only \(\bf \pi r^2\)

OpenStudy (anonymous):

so instead of 10 is it 5?

OpenStudy (anonymous):

or is all of it wrong now?

OpenStudy (jdoe0001):

nope, gimme one sec

OpenStudy (jdoe0001):

\(\bf r = 5\qquad \qquad Area = \pi r^2 \implies Area = 25 \pi\\ \textit{area of the other circle is double that } 50\pi \textit{ thus }\\ 50 \pi = \pi r^2 \implies \sqrt{\cfrac{50\cancel{\pi}}{\cancel{\pi}} }= r \implies r = \sqrt{50}\)

OpenStudy (jdoe0001):

"translated to the right 4 units and up 2 units" => \(\bf (x+4)^2+(y-2)^2 = r^2 \implies (x+4)^2+(y-2)^2 = 50\)

OpenStudy (anonymous):

got it thanks

OpenStudy (jdoe0001):

yw